# DISPATCH - Decentralized Informed Spatial Planning and Assignment of Tasks for Cooperative Heterogeneous Agents

### v3

@@ Abstract
We evaluate on Multi-Agent Particle Environment (MPE) simulations across varying team sizes against centralized EG, Hungarian, and Min–Max distance baselines, and also present a Webots-based warehouse proof-of-concept with heterogeneous robots. Both methods preserve the fairness–efficiency balance of the EG solution under partial observability, with EG-MARL achieving near-centralized coordination and reduced travel distances, and the online mechanism enabling real-time allocation with competitive fairness.
@@ I Introduction
Teams of autonomous agents are increasingly deployed to tackle missions where tasks are spread across space and subject to strict resource constraints. Such domains include disaster response (e.g., search and rescue) [12, 9], swarm environmental monitoring [3, 6], and planetary exploration [5]. In these applications, agents must navigate through complex environments and perform service operations at designated sites. Tasks often require diverse capabilities or skill levels—for instance, transport, delicate manipulation, or specialized inspection—making assignment both a problem of navigation and matching heterogeneous agents to compatible service demands. These challenges are further compounded by travel and service costs, limited energy resources, and environmental uncertainty [11]. This motivates the need for principled and adaptive approaches to multi-agent task allocation.
A key challenge in multi-agent task allocation is balancing efficiency with fairness across tasks. Prior approaches in robotics and multi-agent planning—such as combinatorial auctions and deep MARL frameworks—typically aim to maximize collective performance [8]. However, methods that solely optimize overall system efficiency often create inequities: some tasks are completed quickly and matched to highly capable agents, while others face long delays or poor matches [26, 27]. In mission-critical or human-facing domains, such disparities can undermine reliability and erode trust in autonomous systems. This gap is especially pronounced in spatial, dynamic, and online settings, where tasks differ in urgency, workload, and compatibility with agents.
To address these fairness limitations, classical economic models offer a complementary perspective grounded in equilibrium theory. Competitive equilibria and Eisenberg–Gale (EG) programs provide principled mechanisms that balance efficiency and fairness by guaranteeing Pareto-optimal and envy-free outcomes under weighted budgets [4, 7]. However, these formulations traditionally assume centralized and static settings, and rarely account for embodied agents navigating dynamic, partially observable environments.
We study a multi-agent navigation-and-service problem in a two-dimensional workspace with static obstacles, where $n$ heterogeneous agents must reach and serve $m$ spatially distributed tasks under partial observability. Each task has a workload demand and priority, and agents differ in service capabilities. The preference vector encodes task–agent compatibility. In addition, the tasks also have a budget/weight that signifies their importance. The objective is to minimize total completion time and travel cost while ensuring a fair workload distribution across tasks with respect to preferences and the importance of tasks.
@@
To address these challenges, our work bridges a gap between equilibrium theory and decentralized multi-agent coordination. We extend the Eisenberg–Gale framework beyond its classical static and centralized setting to embodied agents with spatial costs and partial observability, and show how its fairness structure can be used to guide decentralized policy learning in heterogeneous teams. This coupling unifies market-based fairness and motivates two complementary algorithmic realizations. We embed a spatially modified Eisenberg–Gale convex program as a centralized teacher within a CTDE MARL framework to learn decentralized, fairness-aware policies, and we pair this with an EG-inspired online assignment method that provides a centralized upper-bound baseline. This contrasts with prior market‑based approaches such as FMC‑TA [10], FMC_TA+ [23], and online Fisher/EG pricing [15], which assume centralized global information, and with prior fair‑MARL methods [13, 14, 2], which rely on reward shaping or max–min objectives without an equilibrium‑based fairness–efficiency reference.
Fairness Notions. We adopt two complementary notions of fairness. First, the centralized Eisenberg–Gale (EG) program provides a Pareto-efficient benchmark: no task’s weighted utility can increase without decreasing another’s. Because EG-MARL operates under partial observability, we measure how closely it approximates this benchmark through regret, the gap between its realized EG objective and the centralized optimum. Second, we assess distributional fairness using the normalized utility-to-weight ratio $\rho_{j}=u_{j}^{\text{realized}}/w_{j}$, with uniformity quantified by $F(\rho)=1/\mathrm{CV}(\rho)$. This captures how evenly service is delivered across tasks relative to their importance.
To provide a complementary perspective on distributional balance, we also report Jain’s Fairness Index,
@@
which equals 1 when all tasks receive identical normalized utility and decreases as disparities grow. As depicted in Fig. 1, our setting requires both spatially aware navigation and skill-aligned assignment, where skill compatibility is modeled through the preference coefficients. Our main contributions are summarized as follows:
@@
We validate our approach in Multi-Agent Particle Environment (MPE) simulations across multiple team sizes and in a Webots warehouse proof-of-concept with heterogeneous robots.
@@ II Background and Related Work
Fairness in multi-agent systems was initially studied under planner-driven optimization, where a single controller allocates resources to maximize social welfare [21, 33]. Lan and Chiang [18] established an axiomatic foundation for fairness through $\alpha$-fairness and proportional fairness, framing it as a welfare-maximization problem. Market-based formulations such as the Fisher market [35] and the Eisenberg–Gale (EG) program [34] later introduced equilibrium mechanisms that guarantee Pareto efficiency and envy-freeness under budget constraints [7, 4], but these frameworks assumed complete global information and static environments. To address temporal and spatial heterogeneity, Amador et al. [10] proposed a Fisher-market-inspired algorithm (FMC-TA) that produces dynamic Pareto-efficient assignments incorporating travel costs. Building on this foundation, Amador and Zivan [23] introduced FMC_TA+, which extends the market-clearing formulation to concave personal utilities and heterogeneous agent skills. Their results showed that slight concavity encourages cooperative task execution among diverse agents while maintaining Pareto-optimal and envy-free equilibria. More recently, Jalota and Ye [15] generalized EG and Fisher markets to stochastic and online environments through adaptive pricing updates, ensuring near-optimal efficiency and bounded regret. These advances mark a shift from static, centralized equilibria toward dynamic and sequential decision-making, where fairness must emerge through local interactions and learning rather than global coordination.
Extending competitive equilibrium formulations with multi-agent reinforcement learning(MARL) offers a natural path forward. A foundational direction by Zhang and Shah [13] introduced a regularized max–min fairness criterion for factored multi-agent MDPs, maximizing the minimum long-term reward while maintaining Pareto efficiency. Their framework established the theoretical link between fairness and efficiency, but assumes centralized control and full observability. Jiang and Lu [14] extended this to decentralized settings through the Fair-Efficient Network (FEN), which learns joint fairness–efficiency tradeoffs using hierarchical controllers and local consensus. Their approach demonstrated that fairness principles can be embedded directly into decentralized policy learning.
Beyond centralized formulations, many works embed fairness via reward shaping, which encodes cooperative objectives or fairness criteria into the learning signal to improve stability and coordination in multi-agent systems [26, 27]. In fairness-aware settings, such techniques modify the agents’ rewards to penalize disparities or encourage balanced contributions to shared outcomes. Aloor et al. [2] further demonstrated how fairness can be enforced through distance-based reward shaping, enabling cooperative agents to maintain equitable service distribution in spatial coordination tasks.
@@ III Preliminaries
CTDE enables agents to train using global information during training but operate with local observations at execution time, and has been widely discussed in recent surveys [25]. Our CTDE formulation builds on the design principles introduced in InforMARL [1], which demonstrated that local graph-based message passing can enable decentralized policies to scale under partial observability.
@@ IV EG-MARL Framework
$x^{(i)}_{j}=[\,\mathbf{v}_{i}^{j},\;\mathbf{p}_{i}^{j},\;\mathbf{p}_{i}^{\text{goal},j},\;a_{j},\;t_{j},\;\text{pref}_{ji}^{\text{task}},;w_{j},\;w_{j},\;\mathbf{c}_{j},\;e_{j}\,],$
@@
Multi-robot deployments for exploration, monitoring, and search-and-rescue frequently operate under communication limits due to attenuation, interference, and obstructions. As a result, reliable communication is typically restricted to local neighborhoods, and coordination must depend on direct or multi-hop message passing.

### v4

@@ Abstract
We evaluate on Multi-Agent Particle Environment (MPE) simulations across varying team sizes against centralized EG, Hungarian, and Min–Max distance baselines, and also present a Webots-based warehouse proof-of-concept with heterogeneous robots. Both methods preserve the fairness–efficiency balance of the EG solution under partial observability, with EG-MARL achieving near-centralized coordination and reduced travel distances, and the online mechanism enabling real-time allocation with competitive fairness. Together, these results demonstrate that spatially aware EG formulations can effectively guide decentralized coordination in agents with heterogeneous capabilities.
@@ I Introduction
Teams of autonomous agents are increasingly deployed to tackle missions where tasks are spread across space and subject to strict resource constraints. Such domains include disaster response (e.g., search and rescue) [12], swarm environmental monitoring [3, 6], and planetary exploration [5]. In these applications, agents must navigate through complex environments and perform service operations at designated sites. Tasks often require diverse capabilities or skill levels—for instance, transport, delicate manipulation, or specialized inspection—making assignment both a problem of navigation and matching heterogeneous agents to compatible service demands. These challenges are further compounded by travel and service costs, limited energy resources, and environmental uncertainty [11].
A key challenge in multi-agent task allocation is balancing efficiency with fairness across tasks. Prior approaches in robotics and multi-agent planning aim to maximize collective performance [8]. However, methods that solely optimize overall system efficiency often create inequities: some tasks are completed quickly and matched to highly capable agents, while others face long delays or poor matches. Chin et al. [9] studied drone-base placement for post-disaster search and rescue and show that purely efficiency-driven deployment can create substantial inequities in access across affected regions. In mission-critical or human-facing domains, such disparities can undermine reliability and erode trust in autonomous systems. Imposing fairness constraints can reduce system efficiency, highlighting the fundamental fairness–efficiency trade-off in multi-agent decision-making [27]. This is especially pronounced in spatial, dynamic, and online settings, where tasks differ in urgency, workload, and compatibility with agents.
We study a multi-agent navigation-and-service problem in a two-dimensional workspace with static obstacles, where $n$ heterogeneous agents must reach and serve $m$ spatially distributed tasks under partial observability. Tasks have workload demands and importance weights, while agents differ in service capabilities, captured through a preference vector encoding agent–task compatibility. The goal is to minimize total completion time and travel cost while ensuring fair service across tasks given their preferences and importance.
@@
To address these challenges, we connect equilibrium theory with decentralized multi-agent coordination. We extend the Eisenberg–Gale framework to embodied agents with spatial costs and partial observability, showing how its fairness structure can guide decentralized policy learning. This yields two complementary algorithms: (i) a CTDE MARL framework that uses a spatially modified EG program as a centralized teacher, and (ii) an EG-inspired online assignment method that provides a centralized baseline. This differs from prior market-based methods such as FMC-TA [10], FMC_TA+ [20], and online Fisher/EG pricing [15], which assume centralized global information, as well as from fair-MARL approaches [13, 14, 2], which rely on reward shaping or max–min objectives without an equilibrium-based fairness–efficiency reference.
Fairness Notions. We adopt two complementary notions of fairness. First, the centralized Eisenberg–Gale (EG) program provides a Pareto-efficient benchmark. Because EG-MARL operates under partial observability, we measure deviation from this benchmark using regret, defined as the gap between its realized EG objective and the centralized optimum.
Second, we evaluate distributional fairness via the normalized utility-to-weight ratio $\rho_{j}=u_{j}^{\text{realized}}/w_{j}$, with uniformity quantified by $F(\rho)=1/\mathrm{CV}(\rho)$. We additionally report Jain’s Fairness Index,
@@
which equals 1 under perfectly equal service and decreases as disparities increase. Our setting requires both spatially aware navigation and skill-aligned assignment, where skill compatibility is modeled through the preference coefficients. Our main contributions are summarized as follows:
@@
We validate our approach in Multi-Agent Particle Environment (MPE) simulations across multiple team sizes and in a Webots warehouse (See Appendix), proof-of-concept with heterogeneous robots.
@@ II Background and Related Work
In multi-agent settings, fairness was studied under planner-driven optimization, where a single controller allocates resources to maximize social welfare [19]. Lan and Chiang [17] established an axiomatic foundation for fairness through $\alpha$-fairness and proportional fairness, framing it as a welfare-maximization problem. Market-based formulations such as the Fisher market [29] and the Eisenberg–Gale (EG) program [28] later introduced equilibrium mechanisms that guarantee Pareto efficiency and envy-freeness under budget constraints [7, 4]. More recently, Kumar and Yeoh proposed DECAF [18], which uses DDQN to learn centralized allocation policies that trade off total utility and long-term fairness under multiple fairness objectives. To address temporal and spatial heterogeneity, Amador et al. [10] proposed a Fisher-market-inspired algorithm (FMC-TA) that produces dynamic Pareto-efficient assignments incorporating travel costs. Building on this foundation, Amador and Zivan [20] introduced FMC_TA+, which extends the market-clearing formulation to concave personal utilities and heterogeneous agent skills. Their results showed that slight concavity encourages cooperative task execution among diverse agents while maintaining Pareto-optimal and envy-free equilibria. Zhang and Shah [13] introduced a regularized max–min fairness criterion for factored multi-agent MDPs, linking fairness and efficiency, though their formulation assumes centralized control and full observability.However, the aforementioned frameworks assumed complete global information. More recently, Jalota and Ye [15] generalized EG and Fisher markets to stochastic and online environments through adaptive pricing updates, ensuring near-optimal efficiency and bounded regret. Zimmer et al. [24] propose a decentralized cooperative MARL framework that balances efficiency and equity by optimizing a social welfare function, enabling inequality-sensitive objectives such as max–min and generalized Gini–based fairness. These advances mark a shift from static, centralized equilibria toward dynamic and sequential decision-making, where fairness must emerge through local interactions and learning rather than global coordination.
Complementary to these equilibrium-inspired methods, many works introduce fairness through reward shaping, adjusting agents’ rewards to penalize disparities or promote balanced contributions [21]. Jiang and Lu [14] proposed a decentralized hierarchical Multi-Agent Reinforcement Learning method that encourages equitable outcomes by optimizing a shaped “fair-efficient” reward combining system utility with a penalty for agents’ deviation from the population-average utility. Aloor et al. [2] further showed that distance-based shaping can enforce equitable service distribution in spatial coordination tasks.
@@ III Preliminaries
CTDE enables agents to train using global information during training but operate with local observations at execution time, and has been widely discussed in recent surveys [22, 23]. Our CTDE formulation builds on the design principles introduced in InforMARL [1], which demonstrated that local graph-based message passing can enable decentralized policies to scale under partial observability.
@@ IV EG-MARL Framework
$x^{(i)}_{j}=[\,\mathbf{v}_{i}^{j},\;\mathbf{p}_{i}^{j},\;\mathbf{p}_{i}^{\text{goal},j},\;a_{j},\;t_{j},\;\text{pref}_{ji}^{\text{task}},\;w_{j},\;\mathbf{c}_{j},\;e_{j}\,],$
@@
Multi-robot deployments for exploration, monitoring, and search-and-rescue frequently operate under communication limits due to attenuation, interference, and obstructions [26]. As a result, reliable communication is typically restricted to local neighborhoods, and coordination must depend on direct or multi-hop message passing.